Calculate the current ratio for each business below. Which business has the best short term solvency position given your calculations? SHOW WORK

Company Current Assets Total Assets Current Liabilities Total Liabilities Net Income Current Ratio
Alpha 74,524 168,672 60,100 150,600 94,958
Beta 207,536 290,290 152,600 203,000 207,536
Gamma 60,125 66,929 32,500 52,700 36,725
Delta 95,335 182,520 82,900 135,200 105,283
Answer #1

Alpha

Current ratio = Current Assets / Current liabilities

= 74,524 / 60,100

= 1.24

Beta

Current ratio = Current Assets / Current liabilities

= 207,536 / 152,600

= 1.36

Gamma

Current ratio = Current Assets / Current liabilities

= 60,125 / 32,500

= 1.85

Delta

Current ratio = Current Assets / Current liabilities

= 95,335 / 82,900

= 1.15

Gamma has the best short term solvency position as its current ratio is 1.85